OpenStudy (anonymous):

A bag of cement mix has a mass of 15.6 kilograms. Which number sentence will tell how many grams of mass the bag of cement mix has? A. 15.6 × 1,000 B. 15.6 ÷ 100 C. 15.6 ÷ 1,000 D. 15.6 × 100

OpenStudy (anonymous):

@Tgannon998

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

I don't know

OpenStudy (tgannon998):

How many grams in a kilogram

OpenStudy (anonymous):

10000

OpenStudy (anonymous):

1000

OpenStudy (tgannon998):

1000 grams = 1 kilogram

OpenStudy (anonymous):

ok

OpenStudy (tgannon998):

so you would have to multiply 15.6 by 1000

OpenStudy (anonymous):

15600

OpenStudy (tgannon998):

yes so the answer is A

OpenStudy (anonymous):

@flower2468

OpenStudy (flower2468):

I think they are right about the answer

OpenStudy (anonymous):

An average bathtub holds about 110,000 milliliters of water. How many liters of water does the average bathtub hold?

OpenStudy (flower2468):

1milliliter=0.001 liter so 110,000*0.001=......?

OpenStudy (tgannon998):

or 110,000/1000

OpenStudy (anonymous):

110

OpenStudy (tgannon998):

yes

OpenStudy (anonymous):

Olga’s family traveled 942 kilometers on the first day of their trip and 894.5 kilometers on the second day. How many more meters did Olga’s family travel on the first day?

OpenStudy (flower2468):

1kilometer=100meters so 942*100=.....? and 894.5*100=....?

OpenStudy (tgannon998):

subtract the second number from the first number

OpenStudy (tgannon998):

then multiply it by 1000

OpenStudy (anonymous):

89540

OpenStudy (anonymous):

i got hte answer

OpenStudy (anonymous):

HELLO?

OpenStudy (flower2468):

you got the first one what about the second one?

OpenStudy (anonymous):

i don't know

OpenStudy (tgannon998):

you final answer should be 47500 right?

OpenStudy (anonymous):

yes

OpenStudy (tgannon998):

medal?

OpenStudy (anonymous):

The oil company removed 5,421 liters of gas from a tank that contained 5,855 liters. How many milliliters of gas are left in the tank?

OpenStudy (anonymous):

@flower2468

OpenStudy (tgannon998):

434000

OpenStudy (anonymous):

is the answer?

OpenStudy (tgannon998):

yes